You flip 3 coins. Let X, the random variable, be the number of tails on all three coins. Construct a probability distribution for this experiment.



1
Expert's answer
2022-02-02T09:40:04-0500

Solution:

S = {HHH, HHT, HTH, THH, HTT, THT, TTH, TTT}

Total outcomes = 8

X, the number of tails on all three coins = {0,1,2,3}

P(0)=1/8 (one out of eight outcomes,i.e., HHH)

P(X=1)=3/8 (three out of eight outcomes,i.e., HHT, HTH, THH)

P(X=2)=3/8 (three out of eight outcomes,i.e., HTT, THT, TTH)

P(X=3)=1/8 (one out of eight outcomes,i.e., TTT)